OpenStudy (anonymous):

is anyone good with Inverse of exponential functions? If so can someone help me with a few questions? I will fan and medal.

OpenStudy (anonymous):

Given f(x)=3^x, find f^–1(x) and f^–1(27).

OpenStudy (anonymous):

@zepdrix help please?

OpenStudy (kainui):

The inverse of an exponential function is a logarithm, they're sorta weird if you're new to them, so here's an example: \[ \log_2(8)\] The little 2 on the log is the base, and this is basically saying, "what power of 2 is this?" So since \(8=2^3\) then we can write: \[ \log_2(8)=3\] so the trick is if we have something like \(y=2^x\) the inverse will be to take the logarithm base 2 of both sides since: \[ \log_2(y) = \log_2(2^x)\] Notice that right side is saying what's the power on \(2^x\) ? Well that's just x! So we rewrite it: \[ \log_2(y) = x\] Since this is the inverse function we switch y and x (or in your case f(x) and x. to get: \[ \log_2(x) = y\]

OpenStudy (anonymous):

Okay well that kinda makes sense

OpenStudy (anonymous):

but can you like sorta walk me through this one? Given f(x)=3^x, find f^–1(x) and f^–1(27). @Kainui

OpenStudy (kainui):

Well I walked you through some stuff right there, see if you can take what I'm saying and apply it to solve your problem first and I'll help you.